# RED-DiffEq: Regularization by denoising diffusion models for solving inverse PDE problems with application to full waveform inversion

### v1

@@ Abstract
Partial differential equation (PDE)-governed inverse problems are fundamental across various scientific and engineering applications; yet they face significant challenges due to nonlinearity, ill-posedness, and sensitivity to noise. Here, we introduce a new computational framework, RED-DiffEq, by integrating physics-driven inversion and data-driven learning. RED-DiffEq leverages pretrained diffusion models as a regularization mechanism for PDE-governed inverse problems. We apply RED-DiffEq to solve the full waveform inversion problem in geophysics, a challenging seismic imaging technique that seeks to reconstruct high-resolution subsurface velocity models from seismic measurement data. Our method shows enhanced accuracy and robustness compared to conventional methods. Additionally, it exhibits strong generalization ability to more complex velocity models that the diffusion model is not trained on. Our framework can also be directly applied to diverse PDE-governed inverse problems.
@@ 1 Introduction
Inverse problems governed by partial differential equations (PDEs) play a crucial role in diverse scientific and engineering domains [1, 2, 3, 4], from medical imaging [5] and fluid dynamics [6] to geophysical studies [7]. These problems involve inferring unknown parameters or fields from indirect observations, where the forward process is described by PDEs. A prominent example is full waveform inversion (FWI) in seismic imaging, which aims to infer subsurface structures for resource exploration, environmental studies, and seismic hazard assessment [8]. Unlike linear inverse problems, PDE-governed inverse problems usually formulate the inference as a nonconvex and nonlinear optimization task, iteratively minimizing the mismatch between observed data and current PDE solution to optimize unknown parameters [9].
These inverse problems are intrinsically challenging due to their nonlinearity and ill-posedness, making them highly sensitive to stochastic perturbations, such as measurement noise and missing data [9, 10, 11, 12]. In FWI, for instance, these challenges manifest as cycle skipping and convergence to local minima, particularly in complex settings or with limited data quality [13]. Various regularization strategies have been proposed to address these issues. Classical approaches, such as Tikhonov regularization, promote solution smoothness [14], while total variation (TV) regularization preserves sharp boundaries [15, 16]. These regularization techniques enhance the optimization process, improve convergence, and mitigate the effects of ill-posedness. They remain state-of-the-art regularization methods in full waveform inversion [17].
@@
Machine learning has introduced innovative approaches to PDE-governed inverse problems. Deep learning models, particularly convolutional neural networks, have demonstrated success in learning direct mappings from observations to underlying parameters in synthetic environments [18, 19, 20]. These approaches offer potential advantages in computational efficiency and eliminate sensitivity to initial parameters [21]. However, their practical application is often limited by poor generalization, especially when handling noise, missing data, or scenarios not represented in training datasets [22, 23, 24]. To address these limitations, advanced physics-informed machine learning frameworks have emerged. Physics-informed neural networks (PINNs) integrate physical constraints directly into the learning process, demonstrating promising results for solving inverse problems [25, 26, 27, 28, 29, 30, 31, 32]. This integration enhances solution consistency with known physical laws while maintaining computational efficiency [33, 34, 35, 36]. Building on this foundation, deep neural operators [37, 38, 39, 40, 41] have advanced the field by learning sophisticated mappings between function spaces, offering improved generalization across diverse physical scenarios [42, 43, 44, 45, 46, 47]. Despite these significant advances, both approaches continue to face challenges with out-of-distribution data, particularly when confronted with complex real-world problems.
@@
Another promising recent development in machine learning has been the emergence of diffusion models, which have revolutionized generative modeling through their iterative denoising approach to data generation, effectively capturing complex probability distributions [48, 49, 50]. These models offer unique advantages for inverse problems due to their ability to incorporate prior information and their inherent robustness to noise [51, 52, 53, 54]. Diffusion models have also demonstrated success in reconstructing complex fields from limited or noisy observations in various domains [55, 56, 57]. However, their application to inverse problems involving explicit PDE solvers as forward operators remains under-explored. In FWI, for instance, initial efforts have primarily focused on leveraging diffusion models as prior regularizers [58, 59, 60], or leveraging a joint diffusion architecture to solve the velocity model in the latent space [61]. These approaches, while promising in synthetic scenarios, have not yet always outperformed conventional regularization techniques like Tikhonov regularization.
@@
Here, we develop a general framework, regularization by denoising using diffusion models for partial differential equations (RED-DiffEq), that integrates diffusion models directly into PDE-governed inverse problems as a regularization mechanism. Our approach leverages diffusion models to learn robust prior distributions over plausible solutions from synthetic datasets. By employing the pretrained diffusion model as a physics-aware regularization term, the regularization effect is achieved by calculating the residual between actual and predicted noise.
@@
We demonstrate the effectiveness of RED-DiffEq through extensive validation on FWI, a challenging inverse problem in geophysics. Our results consistently show superior performance in accuracy and robustness compared to conventional regularization methods and existing diffusion-based approaches, both quantitatively and qualitatively. Notably, when trained on the OpenFWI dataset [62], RED-DiffEq demonstrates strong generalization ability to benchmark models such as Marmousi [63] and Overthrust [64]. This highlights its potential for application to real-world field data.
@@ 2 Results
We first introduce full waveform inversion in Section 2.1 and present the RED-DiffEq framework in Section 2.2. We then evaluate the effectiveness of RED-DiffEq in three scenarios from the OpenFWI benchmark (Section 2.3). We also test RED-DiffEq on the Marmousi (Section 2.4) and Overthrust (Section 2.5) data to demonstrate its generalizability.
@@ 2.1 Full waveform inversion
Full waveform inversion seeks to reconstruct high-resolution subsurface velocity models from seismic measurement data (potentially missing or noisy) by leveraging the physics of wave propagation (Fig. 1). In the acoustic approximation, wave propagation is governed by
$\frac{1}{\mathbf{x}^{2}(\mathbf{r})}\frac{\partial^{2}\mathbf{u}(\mathbf{r},t)}{\partial t^{2}}\;-\;\nabla^{2}\mathbf{u}(\mathbf{r},t)\;=\;q(\mathbf{r},t),$
@@
Traditional physics-driven numerical methods for FWI are formulated as an optimization problem that iteratively adjusts the velocity model x to minimize an objective function, which reconciles simulated data with observed data:
$\underset{\mathbf{x}}{\text{argmin}}\ \|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|^{2}_{2}+\lambda\cdot R(\mathbf{x}).$ (1)
Here, the first term measures the misfit between the observed seismic data, $\mathbf{u}_{\text{data}}$, and the simulated data, $f_{\text{PDE}}(\mathbf{x})$. The second term is a regularization term weighted by $\lambda$, which is crucial for constraining the solution to be geologically plausible and for improving the convergence of the inversion.
While traditional regularization methods, such as Tikhonov and TV, improve performance, they often fail to produce geologically realistic models. Tikhonov regularization, for instance, tends to introduce over-smoothing, while TV can create “staircase” artifacts. Moreover, this physics-driven approach leads to an ill-posed inverse problem that is highly sensitive to the choice of the initial velocity model and the presence of measurement noise. For a more detailed description of FWI, please refer to Section S1.
Figure 1: Forward modeling and full waveform inversion. The velocity model (top) with five sources (A–E) generates seismic data under different conditions: clean data (noise-free), Gaussian noise, and missing traces.
@@ 2.2 RED-DiffEq for full waveform inversion
To solve inverse problems of PDEs like FWI, we develop a new regularization framework named RED-DiffEq (Algorithm S1) based on the regularization by denoising (RED) method [65]. Let $p_{\text{data}}$ denote the density of the prior distribution over clean, geologically plausible velocity models. For a velocity model $\mathbf{x}$ at iteration $k$ (denoted $\mathbf{x}_{k}$), the RED regularizer is
@@
$\mathcal{D}_{\theta}$ is a Tweedie-inspired denoise operator that nudges $\mathbf{x}_{k}$ toward regions of higher density. To define $\mathcal{D}_{\theta}$, let $\gamma(t)\!\in\!(0,1]$ be a noise schedule (Algorithm S2), where $t\sim\mathcal{U}\{1{:}T\}$, and $T$ is a hyperparameter. In each iteration of RED-DiffEq, we first compute a noisy velocity model $\mathbf{x}_{k,t}$ from the Variance-Preserving (VP) corruption function by
@@
We approximate this score function by a neural network $\hat{\boldsymbol{\epsilon}}_{\theta}$ (typically a U-Net as in Fig. 2a) using the Denoising Diffusion Probabilistic Model (DDPM) on a corpus of clean velocity models $\mathbf{x}_{\text{data}}$ (Fig. 2b). After training, we estimate the score in Eq. (4) as
@@
Finally, substituting Eq. (6) into Eq. (2) gives the RED regularizer by
$R(\mathbf{x}_{k})\;=\;\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\,\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)\right],\quad w(t)\;=\;\sqrt{\tfrac{1-\gamma(t)}{\gamma(t)}},$
where $\boldsymbol{\epsilon}$ is the random Gaussian noise we draw in Eq. (3). In our experiments we adopt a pragmatic variant that drops $w(t)$ and uses a constant $\lambda$, which empirically improves stability and convergence. Therefore, the empirical inversion objective minimized at iteration $k$ is given by
$\mathcal{L}(\mathbf{x}_{k})\;=\;\bigl\|\,\mathbf{u}_{\text{data}}-f_{\mathrm{PDE}}(\mathbf{x}_{k})\,\bigr\|_{2}^{2}\;+\;\lambda\,\underbrace{\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)}_{\hat{R}(\mathbf{x}_{k})},$
where $\hat{R}$ provides an unbiased Monte Carlo (MC) estimator of the RED regularizer (Fig. 3). In practice, we randomly sample $(t,\boldsymbol{\epsilon})$ at each iteration, which injects stochasticity into the inversion process. We present the details of RED-DiffEq in Sec. 4.1.
Figure 2: Diffusion model architecture and generated samples of velocity map. (a) Schematic illustration of the U-Net architecture with time embedding used in the diffusion model. (b) Overview of the complete diffusion process, showing both the forward noising process and the learned reverse denoising process. (c) Comparison between velocity maps from the training dataset (top row) and unconditionally generated velocity maps from a single pretrained diffusion model (bottom row).
Figure 3: Schematic illustration of RED-DiffEq for full waveform inversion. (a) An overview of the inversion process that iteratively updates the velocity model. (b) An illustration of each iteration step of RED-DiffEq. (c) Calculation of the diffusion-based regularization term. (d) Post-processing refinement step using the pretrained diffusion model to further refine the velocity map after the main inversion process is finished.

### v2

# arXiv:2509.21659v2[cs.LG] 01 Mar 2026

@@ Abstract
Partial differential equation (PDE)-governed inverse problems are fundamental across various scientific and engineering applications; yet they face significant challenges due to nonlinearity, ill-posedness, and sensitivity to noise. Here, we introduce a new computational framework, regularization by denoising using diffusion models for partial differential equations (RED-DiffEq), by integrating physics-driven inversion and data-driven learning. RED-DiffEq leverages pretrained diffusion models as a regularization mechanism for PDE-governed inverse problems. We apply RED-DiffEq to solve the full waveform inversion problem in geophysics, a challenging seismic imaging technique that seeks to reconstruct high-resolution subsurface velocity models from seismic measurement data. Our method shows enhanced accuracy and robustness compared to benchmark methods. Additionally, it exhibits strong generalization and domain decomposition capacity, enabling the inversion of more complex velocity models with larger domains than those used in training the diffusion model. Our framework can also be directly applied to diverse PDE-governed inverse problems.
@@ 1 Introduction
Inverse problems governed by partial differential equations (PDEs) play a crucial role in diverse scientific and engineering domains [11, 53, 17, 35], from medical imaging [3] and fluid dynamics [5] to geophysical studies [45]. These problems involve inferring unknown parameters or fields from indirect observations, where the forward process is governed by PDEs. Unlike linear inverse problems, PDE-governed inverse problems usually approximate the solution as a nonconvex and nonlinear optimization task iteratively minimizing the mismatch between observed data and the current PDE solution to optimize unknown parameters. A prominent example is full waveform inversion (FWI) in seismic imaging, which aims to infer subsurface structures for resource exploration, environmental studies, and seismic hazard assessment [57, 66]. In FWI, seismic waves are generated at the Earth’s surface and recorded after propagating through the subsurface. The goal is to reconstruct spatially varying wave-speed (velocity) fields by matching simulated waveforms to recorded measurements.
Such inverse problems are intrinsically challenging due to their nonlinearity and ill-posedness, rendering them highly sensitive to data imperfections like measurement noise and incomplete observations [67, 63, 10, 25]. In FWI, a distinct and critical challenge is cycle skipping, which arises when the predicted and observed waveforms are misaligned by more than half a period. This misalignment causes the local optimization to match incorrect phases, trapping the solution in a spurious basin of attraction [42]. Consequently, cycle skipping is the primary mechanism driving convergence to incorrect local minima, while measurement noise and missing data further exacerbate the instability and ill-posedness of the inversion. To mitigate ill-posedness and enhance robustness against imperfect observations (e.g., noise, limited data), various regularization strategies have been employed. Classical approaches, such as Tikhonov regularization, enforce smoothness constraints on the velocity models [4], whereas Total Variation (TV) regularization promotes piecewise-constant structures to preserve sharp interfaces [16, 33]. These regularizers stabilize the inversion by penalizing geologically implausible high-frequency artifacts and constraining the null space, often resulting in more robust optimization behavior. As such, Tikhonov and TV regularization serve as standard baselines in full waveform inversion [1].
@@
Machine learning has introduced innovative approaches to PDE-governed inverse problems. Deep learning models, particularly convolutional neural networks, have demonstrated success in learning direct mappings from observations to underlying parameters in synthetic environments [68, 62, 69, 30]. These approaches offer potential advantages in computational efficiency and eliminate sensitivity to initial parameters [9]. However, their practical application is often limited by poor generalization, especially when handling noise, missing data, or scenarios not represented in training datasets [40, 27, 64]. To address these limitations, advanced physics-informed machine learning frameworks have emerged. Physics-informed neural networks (PINNs) integrate physical constraints directly into the learning process, demonstrating promising results for solving inverse problems [28, 37, 36]. This integration improves the consistency of the solution with governing physical laws while preserving computational efficiency [7, 65, 41, 6]. In the geophysical domain, PINNs have been actively explored for both wave-equation modeling and inversion, demonstrating their capacity to embed wave propagation physics directly into data-driven frameworks [48, 20, 19, 21, 49]. Building on this foundation, deep neural operators [34] have advanced the field by learning sophisticated mappings between function spaces, offering improved generalization across diverse physical scenarios [24, 32, 13, 38, 26, 31]. Neural operators have also been studied for full waveform inversion [70, 22]. Despite these significant advances, both approaches continue to face challenges with out-of-distribution data, particularly when confronted with complex real-world problems.
@@
Another promising recent development in machine learning has been the emergence of diffusion models, which have revolutionized generative modeling through their iterative denoising approach to data generation, effectively capturing complex probability distributions [52, 18, 50]. These models offer unique advantages for inverse problems due to their ability to incorporate prior information and their inherent robustness to noise [29, 51, 14, 60]. Diffusion models have also demonstrated success in reconstructing complex fields from limited or noisy observations in various domains. Compared to PINNs and neural operators, diffusion-model-based solvers are inherently less sensitive to out-of-distribution scenarios, and more robust against noise and sparse data. More importantly, they eliminate the need for requiring paired data for training and offer a more flexible framework for incorporating numerical solvers during the sampling process to regularize the result [46, 47, 39]. However, their application to inverse problems involving explicit PDE solvers as forward operators remains under-explored. In FWI, initial efforts have primarily focused on embedding the conventional FWI update step within the sampling process of a diffusion model. For instance, Wang et al. proposed DiffusionFWI that integrates FWI iterations into the generative sampling trajectory to solve acoustic FWI [58]. Taufik et al. extended this to elastic FWI [55] and subsequently introduced Iterative Latent Variable Refinement (ILVR) to enhance reconstruction quality [8, 54]. Other studies have investigated joint diffusion architectures that solve for velocity models directly in a latent space [59]. While these approaches establish the viability of diffusion models for FWI, they predominantly embed conventional FWI updates within the diffusion sampling trajectory. This formulation may make it difficult to systematically balance data fidelity against the learned prior.
@@
These observations motivate the need for a framework that combines physics-based inversion with strong data-driven priors while retaining scalability and generalization. Here, we develop a general framework, regularization by denoising using diffusion models for partial differential equations (RED-DiffEq), that integrates diffusion models directly into PDE-governed inverse problems as a regularization mechanism. Our approach leverages diffusion models to learn robust prior distributions over plausible solutions from synthetic datasets. By employing the pretrained diffusion model as a physics-aware regularization term, the regularization effect is achieved by calculating the residual between actual and predicted noise. Moreover, RED-DiffEq incorporates a flexible and scalable domain decomposition capability. By utilizing the Domain Decomposed Regularization (DDR) strategy, the framework allows for training on small sub-domains and seamless deployment on substantially larger domains without retraining.
@@
We demonstrate the effectiveness of RED-DiffEq through extensive validation on FWI, a challenging inverse problem in geophysics. Our results consistently show superior performance in accuracy and robustness compared to conventional regularization methods and existing diffusion-based approaches, both quantitatively and qualitatively. Notably, when trained on the OpenFWI dataset (a large-scale synthetic seismic benchmark providing paired velocity models and simulated wavefield recordings [12]), RED-DiffEq demonstrates strong generalization capabilities to benchmark models such as Marmousi [56] and Overthrust [2] enabled by domain decomposition. This highlights its potential for application to real-world field data.
@@ 2 Results
We first introduce full waveform inversion in Section 2.1 and present the RED-DiffEq framework in Section 2.2. We then evaluate the effectiveness of RED-DiffEq in four scenarios from the OpenFWI benchmark (Section 2.3). We also analyze the uncertainty quantification capabilities of our method (Section 2.4). Finally, we test RED-DiffEq on the Marmousi and Overthrust models (Section 2.5) to demonstrate its generalizability and domain decomposition capabilities. All the experimental details, including the training of the diffusion model, hyperparameter tuning, and other design choices are presented in Section S4.
@@ 2.1 Full waveform inversion
Full waveform inversion seeks to reconstruct high-resolution subsurface velocity models from wavefield measurement data (potentially incomplete or noisy) by leveraging the physics of wave propagation (Fig. 1). In the acoustic approximation, wave propagation is governed by
$\frac{1}{\mathbf{x}^{2}(\mathbf{r})}\frac{\partial^{2}\mathbf{u}(\mathbf{r},t)}{\partial t^{2}}\;-\;\nabla^{2}\mathbf{u}(\mathbf{r},t)\;=\;q(\mathbf{r},t),$ (1)
@@
Starting from an initial velocity model $\mathbf{x}$ (typically a smoothed or rough approximation of the true geology), traditional physics-driven numerical methods for FWI are formulated as an optimization problem that iteratively adjusts the velocity model x to minimize an objective function, which reconciles simulated data with observed data:
$\underset{\mathbf{x}}{\text{argmin}}\ \|\mathbf{u}_{\text{data}}-f_{\text{PDE}}(\mathbf{x})\|^{2}_{2}+\lambda R(\mathbf{x}).$ (2)
Here, the first term measures the misfit between the observed seismic data, $\mathbf{u}_{\text{data}}$, and the simulated data, $f_{\text{PDE}}(\mathbf{x})$, obtained from solving Eq. (1). The second term is a regularization term weighted by $\lambda$, which is crucial for constraining the solution to be geologically plausible and for improving the convergence of the inversion.
While traditional regularization methods, such as Tikhonov and TV, improve performance, they often fail to produce geologically realistic models. Tikhonov regularization, for instance, tends to introduce over-smoothing, while TV can create “staircase” artifacts. Moreover, this physics-driven approach leads to an ill-posed inverse problem that is highly sensitive to the choice of regularization method and to the presence of measurement noise that is typically present in seismic measurement data (e.g., Gaussian noise modeling thermal fluctuations or Laplacian noise modeling impulsive outliers with heavy-tailed statistics). For a more detailed description of FWI, please refer to Section S1.
Figure 1: Forward modeling and full waveform inversion. The velocity model (top) with five sources (A–E) generates seismic data under different conditions: clean data (noise-free), noisy data, and data with missing traces.
@@ 2.2 RED-DiffEq for full waveform inversion
To solve inverse problems of PDEs like FWI, we develop a new regularization framework named RED-DiffEq (Algorithm 1) based on the REgularization by Denoising (RED) method [44]. Let $p_{\text{data}}$ denote the probability density of the prior distribution over clean, geologically plausible velocity models. For a velocity model $\mathbf{x}$ at iteration $k$ (denoted $\mathbf{x}_{k}$), the RED regularizer is
@@
$\mathcal{D}_{\theta}$ is a Tweedie-inspired denoising operator that nudges $\mathbf{x}_{k}$ toward regions of higher density. To define $\mathcal{D}_{\theta}$, let $\gamma(t)\!\in\!(0,1]$ be a noise schedule (Algorithm S2), where $t\sim\mathcal{U}\{1{:}T\}$, and $T$ is a hyperparameter. In each iteration of RED-DiffEq, we first compute a noisy velocity model $\mathbf{x}_{k,t}$ from the Variance-Preserving (VP) corruption function by
@@
We approximate this score function by a neural network $\hat{\boldsymbol{\epsilon}}_{\theta}$ (typically a U-Net as in Fig. 2a) using the Denoising Diffusion Probabilistic Model (DDPM) on a corpus of clean velocity models $\mathbf{x}_{\text{data}}$ (Fig. 2b). During training, the network is optimized to predict the specific noise $\boldsymbol{\epsilon}$ added to the input during the VP corruption process. After training, we estimate the score in Eq. (5) as
@@
Finally, substituting Eq. (7) into Eq. (3) gives the RED regularizer as
$R(\mathbf{x}_{k})\;=\;\mathbb{E}_{t,\boldsymbol{\epsilon}}\left[w(t)\,\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)\right],\quad w(t)\;=\;\sqrt{\tfrac{1-\gamma(t)}{\gamma(t)}},$ (8)
where $\boldsymbol{\epsilon}$ is the random Gaussian noise we draw in Eq. (4). In our experiments, we adopt a pragmatic variant that drops $w(t)$ and uses a constant $\lambda$, which empirically improves stability and convergence. Therefore, the empirical inversion objective minimized at iteration $k$ is given by
$\mathcal{L}(\mathbf{x}_{k})\;=\;\bigl\|\,\mathbf{u}_{\text{data}}-f_{\mathrm{PDE}}(\mathbf{x}_{k})\,\bigr\|_{2}^{2}\;+\;\lambda\,\underbrace{\mathbf{x}_{k}^{\top}\!\Bigl(\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{k,t},t)-\boldsymbol{\epsilon}\Bigr)}_{\hat{R}(\mathbf{x}_{k})},$ (9)
where $\hat{R}$ provides a Monte Carlo (MC) estimator of the RED regularizer (Fig. 3). In practice, we randomly sample $(t,\boldsymbol{\epsilon})$ at each iteration, which injects stochasticity into the inversion process. We present the details of RED-DiffEq in Sec. 4.1.
For large-scale inversion domains, we develop the Domain Decomposed Regularization (DDR) strategy (Fig. 3e). Specifically, during the inversion process, the large target domain is partitioned into smaller, overlapping subdomains. The diffusion-based regularization is computed locally on each subdomain using the fixed, pretrained diffusion model, and these overlapping regularizations are subsequently aggregated to form a consistent global regularization. This design effectively decouples the training scale from the inversion scale, allowing a model trained on small domains to be seamlessly deployed on arbitrarily large domains without retraining.
After the inversion process, we optionally apply a post-processing refinement step. Specifically, the inverted velocity model is perturbed to a selected diffusion timestep $t$ using the VP corruption function (Eq. (4)), and then denoised unconditionally using the pretrained diffusion model. This refinement produces visually cleaner results, at the cost of a minor degradation in quantitative metrics (Sec. S5). All the results shown in Sec. 2 are the results without the post-processing refinement step.
Figure 2: Diffusion model architecture and generated samples of velocity map. (a) Schematic of the U-Net denoising network used in the diffusion model. A noisy velocity model is processed through an encoder–decoder U-Net with ResBlocks, downsampling and upsampling layers. Sinusoidal time embeddings are passed through a multilayer perceptron and injected into each ResBlock to condition the network on the diffusion timestep. The network predicts the added noise. (b) Overview of the complete diffusion process, showing both the forward noising process and the learned reverse denoising process. In this study, we set the maximum diffusion time step $T=1000$ (c) Comparison between velocity maps from the training dataset (top row) and unconditionally generated velocity maps from a single diffusion model pretrained with four velocity model families altogether (bottom row).
Figure 3: Schematic illustration of RED-DiffEq for full waveform inversion. (a) An overview of the inversion process that iteratively updates the velocity model. (b) An illustration of each iteration step of RED-DiffEq. (c) Calculation of the diffusion-based regularization term. (d) Optional post-processing refinement step using the pretrained diffusion model to further refine the velocity map after the main inversion process is finished. (e) Domain Decomposed Regularization (DDR) strategy for RED-DiffEq. We can train the diffusion model on a small domain and apply on a larger domain during inversion using sliding windows to compute the regularization.